Use False position method determine the root of the equation

Cosx –xe4=0


1
Expert's answer
2021-04-12T08:22:52-0400
"f(x)=cosx-xe^4"

"\frac {df}{dx}=-sinx-e^4"

"sin x=-e^4"

This function has no extremum points.


"x_i=\frac {af(b)-bf(a)}{f(b)-f(a)}"



answer: 0.0185